# Decentralized Smoothing ADMM for Quantile Regression with Non-Convex Sparse Penalties

### v1

@@ III Smoothing Decentraizled ADMM
Step 2. Consensus among nodes and centralized local minimum: Following [21][Theorem 2], the derivative of the augmented Lagrangian converges toward zero at the rate of $o\left(k^{-\frac{1}{4}}\right)$, as the convergence of variables is on the order of $o\left(k^{-\frac{3}{4}}\right)$ and the growth rate of penalty parameters is on the order of $\Omega\left(k^{\frac{1}{2}}\right)$. Moreover, as the approximation functions converge to the actual functions, we can conclude that the limit points satisfy the KKT conditions. One of the KKT conditions in each node concerns local optimality: $\sum_{j\in\mathcal{N}_{l},j<l}\boldsymbol{\xi}^{*}_{jl}+\sum_{j\in\mathcal{N}_{l},j>l}\boldsymbol{\xi}^{*}_{lj}\in M\partial P_{\lambda,\gamma}\left(\mathbf{w}_{l}^{*}\right)-(\mathbf{X}^{(l)})^{\textbf{T}}\boldsymbol{\Psi}_{l}^{*}$. Moreover, considering the edge equality constraints $\mathbf{g}_{lj}=\mathbf{w}_{l},\quad\mathbf{g}_{jl}=\mathbf{w}_{j},\forall j\in\mathcal{N}_{l},j>l$ implies $\boldsymbol{\xi}^{*}_{lj}\in-\omega\partial_{\mathbf{w}_{l}}{\|\mathbf{w}_{l}^{*}-\mathbf{w}_{j}^{*}\|}$. Thus, there exists $\mathbf{d}_{l}\in-\sum_{j\in\mathcal{N}_{l}}\omega\partial_{\mathbf{w}_{l}}\|\mathbf{w}_{l}^{*}-\mathbf{w}_{j}^{*}\|$, such that $\mathbf{d}_{l}\in M\partial P_{\lambda,\gamma}\left(\mathbf{w}_{l}^{*}\right)-(\mathbf{X}^{(l)})^{\textbf{T}}\boldsymbol{\Psi}_{l}^{*}$. Given the condition for $\omega$ and since each $\partial_{\mathbf{w}_{l}}\|\mathbf{w}_{l}^{*}-\mathbf{w}_{j}^{*}\|=-\partial_{\mathbf{w}_{j}}\|\mathbf{w}_{l}^{*}-\mathbf{w}_{j}^{*}\|$, the consensus among all $\mathbf{w}_{l}^{*}$ is inevitable to satisfy the local optimally in the KKT conditions. Finally, consolidating all parts of the KKT conditions into a centralized form, the limit points also satisfy the centralized KKT conditions. ∎

### v2

@@ III Decentralized Smoothing ADMM
Step 2. Consensus among nodes and centralized local minimum: Following [21][Theorem 2], the derivative of the augmented Lagrangian converges toward zero at the rate of $o\left(k^{-\frac{1}{4}}\right)$, as the convergence of variables is on the order of $o\left(k^{-\frac{3}{4}}\right)$ and the growth rate of penalty parameters is on the order of $\Omega\left(k^{\frac{1}{2}}\right)$. Moreover, as the approximation functions converge to the actual functions, we can conclude that the limit points satisfy the KKT conditions. One of the KKT conditions in each node concerns local optimality: $\sum_{j\in\mathcal{N}_{l},j<l}\boldsymbol{\xi}^{*}_{jl}+\sum_{j\in\mathcal{N}_{l},j>l}\boldsymbol{\xi}^{*}_{lj}\in M\partial P_{\lambda,\gamma}\left(\mathbf{w}_{l}^{*}\right)-(\mathbf{X}^{(l)})^{\textbf{T}}\boldsymbol{\Psi}_{l}^{*}$. Moreover, considering the edge equality constraints $\mathbf{g}_{lj}=\mathbf{w}_{l},\quad\mathbf{g}_{jl}=\mathbf{w}_{j},\forall j\in\mathcal{N}_{l},j>l$ implies $\boldsymbol{\xi}^{*}_{lj}\in-\omega\partial_{\mathbf{w}_{l}}{\|\mathbf{w}_{l}^{*}-\mathbf{w}_{j}^{*}\|_{1}}$. Thus, there exists $\mathbf{d}_{l}\in-\sum_{j\in\mathcal{N}_{l}}\omega\partial_{\mathbf{w}_{l}}\|\mathbf{w}_{l}^{*}-\mathbf{w}_{j}^{*}\|_{1}$, such that $\mathbf{d}_{l}\in M\partial P_{\lambda,\gamma}\left(\mathbf{w}_{l}^{*}\right)-(\mathbf{X}^{(l)})^{\textbf{T}}\boldsymbol{\Psi}_{l}^{*}$. Given the condition for $\omega$ and since each $\partial_{\mathbf{w}_{l}}\|\mathbf{w}_{l}^{*}-\mathbf{w}_{j}^{*}\|_{1}=-\partial_{\mathbf{w}_{j}}\|\mathbf{w}_{l}^{*}-\mathbf{w}_{j}^{*}\|_{1}$, the consensus among all $\mathbf{w}_{l}^{*}$ is inevitable to satisfy the local optimally in the KKT conditions. Finally, consolidating all parts of the KKT conditions into a centralized form, the limit points also satisfy the centralized KKT conditions. ∎